OpenStudy (anonymous):

The vertical asymptote for g(x) = ln(3x 4) + 2 is

OpenStudy (anonymous):

you cant have ln(a) where a < 0

OpenStudy (anonymous):

because the derivative of ln(a) = 1/a, and you can never divide by 0

OpenStudy (anonymous):

so for your equation ln(3x + 4) the 3x + 4 HAS to be greater than 0. ( i presume it is 3x + 4 since i cant see the operator inbetween the x and 4)

OpenStudy (anonymous):

so the asymptote is 3x + 4 = 0 x = -4/3

OpenStudy (anonymous):

thank you